OpenStudy (anonymous):

Find all real solutions of the equation. (Enter your answers as a comma-separated list. If there is no real solution, enter NO REAL SOLUTION.) 1 x − 1 + 1 x + 5 = 8 7

OpenStudy (anonymous):

1/(x-1) + 1/(x+5)= 8/7

OpenStudy (dumbcow):

\[\frac{1}{x-1}+\frac{1}{x+5} = \frac{8}{7}\] \[\frac{x+5}{x+5}*\frac{1}{x-1}+\frac{x-1}{x-1}*\frac{1}{x+5} = \frac{8}{7}\] \[\frac{(x+5)+(x-1)}{(x+5)(x-1)}= \frac{8}{7}\] \[7(2x+4) = 8(x+5)(x-1)\] \[7(2x+4) = 8(x^{2}+4x-5)\] \[14x + 28 = 8x^{2}+32x -40\] \[8x^{2}+18x -68 = 0\] \[4x^{2}+9x-34 =0\] \[4x^{2}-8x+17x-34 = 0\] \[4x(x-2)+17(x-2) =0\] \[(4x+17)(x-2) = 0\] \[x = 2 , x = -17/4\]